# Think Before You Talk: Enhancing Meaningful Dialogue Generation in Full-Duplex Speech Language Models with Planning-Inspired Text Guidance

### v1

@@ 4.1 Experimental Setups
Table 4: Pearson correlation results of diverse turn-taking events with conditional generation of various FD-SLMs evaluated under different temperature configurations.
4.2 Results
We aim to answer two critical Research Questions (RQs) through comprehensive experiments. RQ1: Does our method significantly improve the semantic meaningfulness and coherence of the generated full-duplex dialogue? RQ2: Can our method preserve the natural conversational flow after integrating additional texts into the sequence?
RQ1
To evaluate the semantics of the generated dialogue, the most intuitive method is to conduct human evaluations. However, considering the time, cost, and potential inconsistencies associated with human evaluations, we propose utilizing the GPT-score. Specifically, given a generated double-channel dialogue audio, we first transcribe the audio from each channel into text with sentence-level timestamps. For this process, we employ the stable-ts package (Jian 2023) with the Whisper-large-v3 model. Subsequently, the transcribed sentences from both channels are grouped based on their start times, effectively representing the dialogue between the two speakers in text form along the timeline. Finally, the consolidated text is analyzed using GPT-4o (Hurst et al. 2024), which assesses the meaningfulness and coherence of the dialogue on a scale of 0 to 5 with detailed rubrics. The complete GPT-4o prompts for unconditional and conditional generation settings are shown in Appendix G.
Baseline Comparison. We compare our trained baselines, STI and SCI, with dGSLM in the unconditional generation setting on the Fisher test set (the same setting used in the dGSLM paper). As shown in Table 2, STI and SCI perform significantly better than dGSLM, showing that our baselines are trained reasonably well.
Effectiveness of Our Approach. We then demonstrate the effectiveness of TurnGuide through evaluations in both unconditional and conditional settings. In the ”Conditional - All” scenario, GPT-4o uses ordered transcripts from both channels to assess responses based on user speech. However, we observed that GPT-4o overly prioritizes user-assistant interactions, neglecting the coherence of assistant-generated speech within its channel. To address this, the ”Conditional - Assistant” setting provides only the assistant-channel transcript, enabling a focused analysis of semantic quality within the assistant channel.
We compare TurnGuide with SCI and Moshi TS under different temperatures (as shown in Table 3) and make the following observations. 1) Significant performance improvements over baselines. TurnGuide performs significantly better than SCI and Moshi TS (over 30% performance gain), showing that TurnGuide successfully leverages the text information to boost the meaningfulness and coherence of the generated dialogue. 2) Speech-only methods and text-speech methods excel at different temperatures. SCI consistently performs the best at the largest temperature (1.3), whereas TurnGuide tend to perform the best at temperature 1.0. Additionally, TurnGuide shows more performance gain under lower temperatures, such as 0.8 and 0.9. 3) Focusing more on text generation makes the model perform better. In “TurnGuide L2:1” and “TurnGuide L2:1” settings, we set the training loss weight of text and speech tokens to be 2:1 and 3:1, and the results show that they consistently outperform the standard 1:1 setting (“TurnGuide”). 4) TurnGuide achieves performance comparable to its upper bounds. In Table 3, “TurnGuide GT Text” refers to the GLM-4-Voice model trained with our proposed framework but switches to the ground truth transcriptions from the Fisher dataset. TurnGuide performs similarly with TurnGuide GT Text in the 1:1 loss setting and outperforms in the 2:1 and 3:1 loss settings. Additionally, “TextOnly” represents a model trained exclusively on transcriptions without speech tokens, where TurnGuide delivers close performance (2.98–3.18), highlighting the effectiveness of our approach.
Further validation with perplexity. To facilitate the semantic evaluation using GPT-score, perplexity is employed to further demonstrate the superior semantic abilities of TurnGuide. Specifically, we calculate perplexity using the same ordered transcript used in the GPT-score evaluation. Perplexity is measured with three distinct language models: Fisher-text-only, GLM-4-Voice, and Llama-3.1-8b-Instruct (Dubey et al. 2024). The Fisher-text-only model, trained using Fisher-provided text transcript data on the GLM-4-Voice base model, serves as the ground truth for text-based dialogue distribution. GLM-4-Voice acts as the SLM backbone, while Llama-3.1-8b-Instruct serves as the text-only model backbone. We analyze the perplexity of generated dialogues across different temperature settings in both conditional and unconditional scenarios. As illustrated in Figure 4, TurnGuide consistently achieves lower perplexity compared to baseline approaches, demonstrating superior semantic modeling abilities.
RQ2
To evaluate whether the full-duplex dialogue generated by TurnGuide preserves a natural conversational flow, we measure turn-taking events (Nguyen et al. 2023), categorized into IPU, pause, gap, and overlap, as defined in (Nguyen et al. 2023), with detailed explanations in Appendix H. We calculate the Pearson correlation of the occurrence, cumulative duration, and average duration of the turn-taking events between the 30-second prompts and the corresponding 90-second continuations, as described in (Veluri et al. 2024; Nguyen et al. 2023). This measures whether the model can generate dialogues with similar conversational features to the prompt. The metric is averaged across the four turn-taking events.
Baseline Comparison. We compare our trained baselines, STI and SCI, with dGSLM in the unconditional generation setting (as employed in the dGSLM paper) using the Candor test set. The Candor test set allows for a fairer comparison, addressing the issue where previous papers did not disclose their test set split from the Fisher dataset, leading to significant test set leakage during evaluation. As presented in Table 2, both STI and SCI outperform dGSLM, showing that our baselines are trained reasonably well.
Preservation of the natural conversational flow. We use the same turn-taking evaluation method and compare TurnGuide with SCI and Moshi TS under different temperatures. We evaluate in both unconditional and conditional settings. As shown in Table 4, TurnGuide shows nearly identical performance in Pearson correlation for occurrence, cumulative duration, and averaged duration, showing TurnGuide successfully preserves the natural conversation flow.
5 Conclusion
In this paper, we present a planning-inspired text-guided approach, TurnGuide, to improve the dialogue generation abilities of end-to-end Full-Duplex Speech Language Models. By segmenting assistant speech into turns and training the model with turn-level text-guided speech generation with channel-wise and text-speech interleaving strategies, our approach effectively addresses key challenges of text insertion timing and length, ensuring coherent and natural dialogue generation. Experiments show significant improvements of our method in semantic quality while maintaining natural conversational dynamics, validating its effectiveness for spoken dialogue systems and paving the way for more meaningful and human-like speech interactions.
Appendix A Limitations and Future Works
While the paper introduces TurnGuide, a novel planning-inspired text-guided speech generation approach for enhancing e2e FD-SLMs, it still presents several limitations. First, the effectiveness of TurnGuide is primarily validated on the Fisher and Candor datasets, which, although widely used, may not fully capture the diversity and complexity of real-world conversational scenarios since both Fisher and Candor are telephone conversation datasets. Second, the evaluation relies heavily on GPT-4o-based automated metrics (GPT-score) for semantic assessment, which, despite being aligned with human judgments in most cases, may not perfectly reflect the nuanced human perceptions of dialogue quality.
Figure 5: Illustration of the Moshi training strategy.
Appendix B Judgement Alignment between Human and GPT-4o in Semantic Evaluation
Difference 1 2 3 Overall
Accuracy 7/10 9/10 9/10 25/30
In this paper, we primarily utilize GPT-4o to evaluate the semantics of the generated dialogues. This raises a crucial question: Can AI evaluation scores accurately capture human preferences? To investigate this, we conduct a human listening test to examine the alignment between GPT-evaluated scores and human judgments. For this experiment, we randomly select 30 pairs of generated dialogues that vary in their GPT scores. Specifically, we group these pairs into three categories: 10 pairs with a GPT-score difference of 1, 10 pairs with a difference of 2, and 10 pairs with a difference of 3. Human judges are then asked to evaluate which dialogue in each pair demonstrates better semantics. To ensure unbiased evaluation, the order of dialogues within each pair is randomized. Additionally, judges are provided with both the audio of the dialogues and their transcripts, generated using Whisper-large-v3 with the stable-ts package. We recruit three human judges, all of whom have native English proficiency, to carry out the evaluations.
Appendix C Moshi TS Implementation
First, Moshi utilizes eight codebooks, where every time step includes eight speech tokens for each channel. We adapt this technique for the GLM-4-Voice backbone, which relies on a single codebook method, allotting one speech token per channel per time step. Then, the tokens from each channel are interleaved at each time step, making Moshi rely on a token-level interleaving approach. In the original implementation of Moshi, a text-channel token precedes each assistant-channel speech token, which could either be a genuine text token, a text pad token, or an end-of-text pad token. Moshi also employs text-speech interleaving at token-level, where each text token is incorporated at the precise moment it is spoken. Consequently, if a time step features an audible text token, the corresponding text token is inserted. If not, either a text pad token or an end-of-text pad token is inserted, determined by whether the subsequent text-channel token is an actual text token. We follow the original Moshi implementation by extracting the timestamp of each word with Whisper-timestamped package (Louradour 2023) and the medium Whisper model (Radford et al. 2023). The illustration of our implemented Moshi training strategy is shown in Figure 5.
Appendix D Comparison of Different Training Strategies
@@
We then intuitively explore the reasons behind the lower performance of Moshi TS compared to TurnGuide and identify three key factors contributing to this difference. First, Moshi TS tends to fragment text semantics. Both Moshi TS and our method aim to enhance the semantics of generated speech using textual information. However, Moshi TS splits text tokens so significantly that it hampers the model’s ability to learn coherent text representations. In contrast, TurnGuide aggregates these text tokens, creating a more cohesive grouping that simplifies the learning process for text generation, which leads to notable performance improvements. Second, Moshi TS fails to mitigate the inherently fragmented nature of speech semantics. Here, we propose a new way of understanding the semantic information within a speech audio. Speech audio combines multiple types of information together, including semantic, paralinguistic, and speaker information (Cui et al. 2024), and the semantics within speech can be defined by its transcription, with speech containing semantic information only at the temporal points when each word is sounded. This fragmented nature increases the difficulty for coherent speech generation. Although Moshi TS incorporates text into the speech generation process, it inserts text tokens strictly at the time they are sounded in the speech. This approach only supplements speech with semantic information at points where semantic information already exists, offering limited potential for enhancing semantic representation. By comparison, TurnGuide groups texts within the same turn and processes them collectively. This method introduces denser semantic information, effectively guiding the speech generation process to improve coherence. Third, Moshi TS heavily depends on precise word-level timestamps from the ASR systems. In practice, these timestamps are not always exact, often being slightly ahead or behind the actual sounding moment, which can disrupt text-speech alignment and degrade performance. TurnGuide, however, leverages only the start time of each turn, significantly reducing reliance on word-level timestamp accuracy. This simpler alignment strategy not only lowers the risk of errors but also contributes to stronger overall performance.
Appendix E Fisher and Candor Dataset Split
@@
Both the Fisher and Candor datasets are frequently referenced and employed in e2e FD-SLM studies. However, to date, there is no official training, validation, and test set split for these datasets, nor do existing papers provide details of their dataset splits. This poses a significant challenge for future research, as it becomes difficult to replicate the reported model performance in original studies. Specifically, when researchers use their own dataset splits to train models and subsequently compare their model performance with publicly available checkpoints from prior work, there is a risk of test set leakage. This occurs when an existing model has been trained on test samples included within a new split defined by researchers. To tackle this issue, we propose releasing our train/validation/test splits for the Fisher and Candor datasets and strongly encourage future studies to adopt these splits for a more equitable and reliable basis of comparison. Specifically, Fisher separates all the audio into 117 parts (000-116). We split the train/validation/test set using part IDs 000-110/111-113/114-116, and we randomly select roughly 1,000 samples in our Fisher test set to perform all the evaluations reported in this paper. For the Candor dataset, we randomly select roughly 1,000 samples from the part 001 audio data to perform the evaluations. We will report the IDs of the 1,000 samples for both datasets after the acceptance of the paper.
Appendix F Detailed Computing Infrastructure Specification
We perform all the experiments on A800 GPUs, which have 80GB of GPU memory. The CPU model of the machine is Intel(R) Xeon(R) Platinum 8378A CPU @ 3.00GHz. The operating system is Ubuntu 20.04.4 LTS. Our training environment uses Pytorch 2.0.0 with the transformers library of version 4.49.0. Our inference environment uses Pytorch 1.5.0 with the transformers library of version 4.44.2. We run each experiment once.
@@ Appendix G Semantic Evaluation Prompts
Appendix H Turn-taking Events Explanations
This section details the four turn-taking events utilized in our evaluation: Inter-Pausal Units (IPUs), Pause, Gap, and Overlap. Figure 7 provides examples of these turn-taking events. First, the speech segments within each speaker channel are identified using a Voice Activity Detection (VAD) module. An IPU is defined as continuous speech within a single speaker channel. Multiple speech segments are grouped into a single IPU if the pause between them is shorter than $\tau_{\text{IPU}}$. As previously mentioned, $\tau_{\text{ IPU }}$ is initially set to 0.5 seconds in the multi-modal turn segmentation and alignment framework (Section 3.1). However, for the evaluations, it is adjusted to 0.2 seconds in accordance with prior studies (Nguyen et al. 2023; Veluri et al. 2024; Wang et al. 2025). A Pause refers to the absence of speech within the same speaker channel, while a Gap indicates the discontinuation of speech across different channels. Overlap refers to the time period during which both speakers are speaking simultaneously.

### v2

@@ Appendix B Moshi TS Implementation
Table 7: Pearson correlation results of corpus-level turn-taking events with unconditional and conditional generation of various FD-SLMs evaluated under different temperature configurations.
Appendix C Comparison of Different Training Strategies
@@
We then intuitively explore the reasons behind the lower performance of Moshi TS compared to TurnGuide and identify three key factors contributing to this difference. First, Moshi TS tends to fragment text semantics. Both Moshi TS and our method aim to enhance the semantics of generated speech using textual information. However, Moshi TS splits text tokens so significantly that it hampers the model’s ability to learn coherent text representations. In contrast, TurnGuide aggregates these text tokens, creating a more cohesive grouping that simplifies the learning process for text generation, which leads to notable performance improvements. Second, Moshi TS fails to mitigate the inherently fragmented nature of speech semantics. Here, we propose a new way of understanding the semantic information within a speech audio. Speech audio combines multiple types of information together, including semantic, paralinguistic, and speaker information Cui et al. (2024), and the semantics within speech can be defined by its transcription, with speech containing semantic information only at the temporal points when each word is sounded. This fragmented nature increases the difficulty for coherent speech generation. Although Moshi TS incorporates text into the speech generation process, it inserts text tokens strictly at the time they are sounded in the speech. This approach only supplements speech with semantic information at points where semantic information already exists, offering limited potential for enhancing semantic representation. By comparison, TurnGuide groups texts within the same turn and processes them collectively. This method introduces denser semantic information, effectively guiding the speech generation process to improve coherence. Third, Moshi TS heavily depends on precise word-level timestamps from the ASR systems. In practice, these timestamps are not always exact, often being slightly ahead or behind the actual sounding moment, which can disrupt text-speech alignment and degrade performance. TurnGuide, however, leverages only the start time of each text chunk, significantly reducing reliance on word-level timestamp accuracy. This simpler alignment strategy not only lowers the risk of errors but also contributes to stronger overall performance.
Appendix D Corpus-Level Turn-Taking Evaluations
We also incorporate the results of corpus-level turn-taking evaluations. Following prior works, the turn-taking events considered here include Inter-Pausal Units (IPUs), pause, gap, and overlap (Nguyen et al., 2023). Figure 6 provides examples of these turn-taking events. To give a brief explanation of how these turn-taking events are extracted from the generated dialogue: First, the speech segments within each speaker channel are identified using a pyannote VAD module. An IPU is defined as continuous speech within a single speaker channel. Multiple speech segments are grouped into a single IPU if the pause between them is shorter than $\tau_{\text{IPU}}$. As previously mentioned, $\tau_{\text{ IPU }}$ is initially set to 0.5 seconds in the multi-modal turn segmentation and alignment framework (Section 3.1). However, for the evaluations, it is adjusted to 0.2 seconds in accordance with prior studies Nguyen et al. (2023); Veluri et al. (2024); Wang et al. (2025). A Pause refers to the absence of speech within the same speaker channel, while a Gap indicates the discontinuation of speech across different channels. Overlap refers to the time period during which both speakers are speaking simultaneously.
We calculate the Pearson correlation of the occurrence, cumulative duration, and average duration of the turn-taking events between the 30-second prompts and the corresponding 90-second continuations, as described in Veluri et al. (2024); Nguyen et al. (2023). This measures whether the model can generate dialogues with similar conversational features to the prompt. The metric is averaged across the four turn-taking events.
The corpus-level turn-taking evaluation results on the Fisher testset are presented in Table 7. We conduct evaluations under both unconditional and conditional settings, comparing TurnGuide with various baseline approaches. The results indicate that TurnGuide demonstrates comparable performance to the baselines in terms of Pearson correlation for occurrence, cumulative duration, and average duration. This finding suggests that TurnGuide, which integrates additional text streams into the modeling process, does not exhibit significant behavioral differences at the corpus level.
Remark 2
We include the corpus-level turn-taking results for the sake of completeness. However, as noted in Section 4.2.2, corpus-level statistics are less interpretable and not as suitable for evaluating the turn-taking abilities of FD-SLMs. Therefore, we recommend readers focus on the more reliable results assessed using the Full-Duplex-Bench benchmark (Section 4.2.2).
Appendix E The Fisher Dataset Split
@@
The Fisher dataset is frequently referenced and employed in e2e FD-SLM studies. However, to date, there is no official training, validation, and test set split for the dataset, nor do existing papers provide details of their dataset splits. This poses a significant challenge for future research, as it becomes difficult to replicate the reported model performance in original studies. Specifically, when researchers use their own dataset splits to train models and subsequently compare their model performance with publicly available checkpoints from prior work, there is a risk of test set leakage. This occurs when an existing model has been trained on test samples included within a new split defined by researchers. To tackle this issue, we propose releasing our train/validation/test splits for the Fisher dataset and strongly encourage future studies to adopt these splits for a more equitable and reliable basis of comparison. Specifically, Fisher separates all the audio into 117 parts (000-116). We split the train/validation/test set using part IDs 000-110/111-113/114-116, and we randomly select roughly 1,000 samples in our Fisher test set to perform all the evaluations reported in this paper.
Appendix F Latency Analysis
In this section, we theoretically analyze the overall first-package latency of TurnGuide, which represents the time taken for the user to hear the first output audio sample. Since our model is a fine-tuned version of GLM-4-Voice, much of the latency calculation can refer to the original GLM-4-Voice paper. Specifically, the e2e latency is divided into three components: Speech Tokenization, LLM processing (including pre-filling and token decoding), and Speech Decoding.
With TurnGuide processing speech in chunks, the first-package latency corresponds to the total processing time (Speech Tokenization + LLM + Speech Decoding) for the first chunk. Using a speech chunk size of five tokens, the latency is calculated as follows: 1) The Speech Tokenization time for 5 tokens ($T_{st}$) is 0.01s, the LLM processing time ($T_{LLM}$), which includes pre-filling 5 user speech tokens along with special tokens and decoding a total of 10 assistant tokens (5 text tokens + 5 speech tokens), is 0.3s, and the Speech Decoding time ($T_{SD}$) for 10 speech tokens is 0.75s. Thus, the total latency is $T_{ALL}=T_{st}+T_{LLM}+T_{SD}=1.05s$. It’s worth noting that the optimal speech decoding process involves decoding only 5 speech tokens instead of 10. However, the original GLM-4-Voice vocoder requires at least 10 tokens to decode, meaning the vocoder’s latency can be further reduced in an optimized solution.
@@ Appendix G Semantic Evaluation Prompts
Appendix H Explanations of Turn-Taking Events in Full-Duplex-Bench
This section details the four turn-taking events evaluated in the Full-Duplex-Bench. Specifically, 1) Smooth Turn-taking evaluates the model’s ability to seamlessly take over the conversation after the user’s speaking turn. 2) User Interruption assesses the model’s capability to recognize and adapt to new content introduced by the user during the interaction. 3) Pause Handling determines whether the model can remain silent during brief pauses by the user before they complete their speaking turn. 4) Backchanneling measures the model’s ability to provide brief acknowledgments at appropriate moments while the user is speaking. For a detailed explanations of the evaluation metrics, please refer to their original paper (Lin et al., 2025).
Appendix I Potential Risks
The primary ethical concern regarding TurnGuide involves safety and alignment risks arising from the model’s training data and the absence of post-hoc behavioral tuning. Specifically, since the model is trained on the Fisher dataset—a collection of real-world telephone conversations—it may inherit and replicate biased, sensitive, or harmful topics present in authentic human dialogue. These risks are compounded by the fact that the current iteration has not undergone Reinforcement Learning from Human Feedback (RLHF) (Lambert, 2025; Ouyang et al., 2022; Rafailov et al., 2023) or similar safety alignment procedures to ensure its outputs remain consistent with human preferences. Without a robust refusal mechanism or a safety-oriented backbone, the model’s enhanced capabilities for real-time interruption and overlapping speech could potentially be exploited to generate deceptive or inappropriate content. Consequently, TurnGuide is intended strictly for research purposes and is not suitable for production deployment without the integration of rigorous safety-tuning frameworks and content moderation safeguards.
@@ Appendix I Potential Risks
## Pairwise comparison instructions
### Task overview
You will be shown **30 pairs** of two-speaker dialogue samples generated by different systems.
For each pair, your job is to choose **which dialogue has better semantic quality** (i.e., which one is more meaningful and coherent in what the speakers say).
### What you will receive for each pair
For each pair, you will be given:
- **Audio** of Dialogue A and Dialogue B.
- **Transcripts** for both dialogues (automatically generated with Whisper-large-v3 using the stable-ts package).
The order (A/B) is **randomized** for every pair; do not assume A or B is better.
### What "better semantic quality" means
When deciding which dialogue is better, focus on **content meaning and coherence**, not just fluency.
Prefer the dialogue that is:
- More **understandable** and stays on a recognizable topic.
- More **coherent**, with logical connections between utterances.
- Less **contradictory**, less nonsensical, and less off-topic.
- More **informative/complete** rather than fragmented or empty.
### What to ignore (or de-emphasize)
Do **not** primarily judge based on:
- Audio naturalness (voice quality, prosody, accent, loudness) unless it prevents understanding.
- Turn-taking style (interruptions/overlap) unless it makes the content hard to follow.
- Minor transcription errors, as long as the intended meaning is still clear from audio + context.
### How to make your choice
For each pair, listen to both dialogues (you may replay) and read transcripts as needed.
Then select exactly one option:
- **A is better** (Dialogue A has higher semantic quality).
- **B is better** (Dialogue B has higher semantic quality).
### Consistency and independence
Judge each pair **independently**; do not try to keep scores balanced across the full study.
Make the best decision based only on the evidence in that pair (audio + transcript).
Figure 9: The instructions given to the participants in the pairwise comparison human evaluation.